A box contains 22 red marbles, 52 white marbles, and 10 blue marbles. If a marble is randomly selected from the box, what is the probability that it is red or blue? Express your answer as a simplified fraction or a decimal rounded to four decimal places.
Answer #1

Number of red or blue marbles = 22 + 10 = 32

Number of total marbles = 22 + 52 + 10 = 84

Hence,

P(Red or Blue marble) = 32/84 = 8/21